OpenStudy (anonymous):

1/cosx-sinx=cosx+sinx, solve for all real values for x

OpenStudy (anonymous):

is the cosx-sinx all underneath the 1?

OpenStudy (anonymous):

\[\frac{ 1 }{ \cos x - \sin x }= \cos x + \sin x\]

OpenStudy (anonymous):

cross multiply, and you get the identity